Circular No. 6705

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 O1 A. Beresford, Myrtle Bank, South Australia, reports the visual discov-

ery by Justin Tillbrook of a comet. The following observations are available:

1997 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 22.43  12^{h}28.^{m}4  −18^{o}59′  10 Tillbrook
23.37432 12 20 36.23 −18 09 19.0  Garradd
23.37542 12 20 36.35 −18 09 15.2  _{”}
23.37651 12 20 36.52 −18 09 11.4  _{”}
23.39653 12 20 39.27 −18 08 06.9  _{”}
23.46318 12 20 48.67 −18 04 32.2  _{”}
23.46596 12 20 49.02 −18 04 23.5  _{”}
23.46876 12 20 49.36 −18 04 12.9  _{”}

J. Tillbrook (Clare, S. Australia). 0.20-m reflector. Coma diameter 60′′–

70′′; motion ∼ 3′/hr northward. G. J. Garradd (Loomberah, N.S.W.). 0.25-m f/4.1 reflector + CCD. 30-s

CCD image shows 2′ coma, no tail.

BL LACERTAE J. E. Grove and W. N. Johnson, Naval Research Laboratory, on behalf

of the OSSE team, report: “We have detected emission from BL Lac between 50 and 300 keV with the OSSE instrument on the Compton Gamma Ray Observatory. For the period July 15.7–21.0 UT, the flux in this energy band is (9.0 ± 1.4) × 10^{−4} photons cm^{−2} s^{−1}. There is no evidence for variability. The power-law spectrum is hard, with energy index 0.5 (±0.7, 95-percent confidence). Although OSSE has not previously observed BL Lac, objects in this class are not generally detected in this energy range.”

G. Madejski, Goddard Space Flight Center (GSFC) and University of Maryland; T. Jaffe, GSFC and Hughes/STX; and M. Sikora, Copernicus Center, Warsaw, report the detection of strong x-ray flux from BL Lac, which is undergoing an outburst in the radio, optical, and GeV bands: “The x-ray flux, measured with the Proportinal Counter Array aboard the Rossi X-ray Timing Explorer about every two days between July 16 and 20, increased from ∼ 2 to ∼ 3.6 × 10^{−11} erg s^{−1} cm^{−2} in the band 2–10 keV, with a very hard energy spectrum, well described as a power law with an energy index of 0.4 ± 0.1. This is an increase of flux by about a factor of three, and a very significant hardening (from an energy index of 0.9 ± 0.1) of the x-ray spectrum as compared with the Nov. 1995 ASCA observation.”

1997 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green